# Where is the Truth? The Risk of Getting Confounded in a Continual World

### v2

@@ 5 Conclusion
We have introduced the first continually confounded dataset, ConCon, and have formally described its implications for systematic analysis. In our evaluations, we have empirically demonstrated that even perfect prevention of catastrophic forgetting is insufficient to avoid confounding in disjoint settings, where particularly even joint training fails to generalize. Most importantly, we successfully confirmed the latter to not be true for our strict dataset, where joint training succeeds but continual cumulative training is veritably insufficient due to the manifestation of insidious continual confounding. This stands in stark contrast to the typical assumption that knowledge accumulation is the desirable solution in continual learning. Our work thus provides the means to benchmark and further explore confounding in CL, for example, by investigating causal relationships [31] and related innovative new perspectives on, e.g., causal rehearsal [7] or bias mitigation of spurious correlations [24].
Acknowledgments and Disclosure of Funding
@@
    More training information for the NN (ResNet-18) and NeSy (Neuro-Symbolic concept learner) models.
@@
    Statements on limitations and impact of our paper and the introduced ConCon dataset.
@@
    A datasheet for datasets [13] to document e.g., composition, collection process, and other important aspects.
@@ Appendix A ConCon and the Code Framework
Along with our paper, we also publish the specific ConCon dataset [^3] https://zenodo.org/doi/10.5281/zenodo.10630481, the data generation platform, and the training code to facilitate further research on continual confounding. [^4] https://github.com/ml-research/concon. We include a croissant [1] metadata record in our github repository. [^5] https://github.com/ml-research/concon/blob/main/croissant_concon.json ConCon uses the BSD 2-Clause "Simplified" License, and our code for generating the data and running the experiments is released under the MIT license.
@@ A.1 Dataset Details
Recall that ConCon is built on top of CLEVR [18] and thus consists of images representing geometric objects with different properties. Each image is an RGB image of size 224$\times$224 depicting four objects. In our dataset, an object can be described by four properties, namely shape (cube, sphere, cylinder), material (metal or rubber), size (small or large), and color (gray, red, blue, green, brown, purple, cyan, yellow). Data is generated with respect to the ground truth rule and confounders as described in the main paper. Relationships between objects are not used in ConCon. We show several samples in Figure 4.
As we describe in detail in our main paper, the positive instances in the dataset fulfill the ground truth rule ($g$) and the negative instances do not ($\neg g$). In addition, it is possible that a confounding feature must apply ($c_{t}$) or must not apply ($\neg c_{t}$), depending on the task (denoted by subscript $t$) and the dataset variant (disjoint or strict). The ground truth $g$ is defined as sphere and small cube. The confounders, in the order of tasks, are $c_{1}$: blue, $c_{2}$: metal, and $c_{3}$: large. We summarize the rules that should hold for respective tasks in Tables 2 (disjoint configuration) and 3 (strict configuration).
@@
Table 3: Rules strict: a confounding feature must appear in the positive images but may not appear in the negative images of its respective task. The confounding features may but are not required to appear in images of other tasks.
@@ A.2 Code Framework
It is possible to generate further instances or make changes to the data generation process by using our code repository. The logic in the form of ground truth and confounders, which determine the data generation, can be set in separate files. Note that the intended behavior for our data generation consists of using simple confounders that are only required to be satisfied on a single object. Here, our code could easily be adapted, for example, to support the generation of images based on multi-object confounders by changing the selection of object properties in the respective function. In a similar spirit, both CLEVR and ConCon are generated with the Blender graphics simulation software, allowing for the inclusion of entirely new attributes and confounding features themselves. As such, extensions to, e.g., novel color choices are realizable with either trivial or modest amounts of specification/change in code, or other already available properties (such as relationships) could easily be leveraged in future dataset creation and empirical investigation. Whereas these even more challenging setups may certainly be of interest in the long term, we note that our presently chosen configuration already highlights the catastrophic behavior of investigated methods and exposes insidious continual confounding - making solving our generated ConCon variant first desirable.
@@ Appendix B Experimental Details
Methods. Continual learning methods aim to alleviate forgetting when training and updating models continually. We here give additional information on the CL methods which we consider in our evaluation in the main paper: i) EWC [21] is a method that regularizes weights relevant to the previous task, ii) FROMP [34] stores examples relevant to the previous task and performs functional regularization while training on the current tasks, iii) ER [5] maintains a fixed-size memory buffer of random samples from prior tasks and interleaves them while training on the current task, iv) DER [4] also maintains a memory buffer that stores past samples and predictions, and in addition, the loss function includes a cross-entropy loss for the current task and a regularization term using the stored logits for past tasks, v) BGS [25] tackles the problem of spurious correlations in continual learning by storing samples that are balanced across the class labels and spurious features, and vi) GDUMB [35] stores uniformly balanced past samples but is not a conventional continual learner as the model is trained from scratch at every new task.
@@
NN training. We use a ResNet-18 [16], a convolution neural network-based architecture with skip connections. Following standard deep learning practice, the model is trained with an Adam optimizer [20] with a mini-batch size of 64 and a learning rate of 0.001. The model was trained for 50 epochs for all methods but FROMP, where we trained the NN model for 30 epochs as the performance became worse when training for 50.
@@
NeSy training. We use a neuro-symbolic concept learner (NeSy) [42] that is equipped with a Slot attention model and a Set Transformer. The Slot-Attention module of NeSy is pre-trained on the original CLEVR dataset [27] for the purpose of extracting accurate neuro-symbolic intermediate representations (e.g., disentangled slots for color and shape). We use the (perfectly accurate as a consequence of simulation) structured information from the embeddings of the attention module to then learn only the Set Transformer for the ConCon dataset. The latter is trained using an Adam optimizer [20] with a mini-batch size of 64 and a learning rate of 0.001. The model was trained for 150 epochs. We apply transformations to resize the images to 128$\times$128 to fit the constraints set by prior works [42].
@@
Hyperparameter tuning. We did not excessively tune hyperparameters as all models achieved good solutions when expected in their baseline scenario. As explained in the main body and discussed in much more detail in the next section, observed failure cases are systematic and structural in nature - a consequence of continual learning set-up rather than hyper-parameter values.
@@
Technical Details. We created the dataset using an AMD EPYC 7742 64-Core Processor CPU. All models were built using pytorch and trained and evaluated on an NVIDIA Tesla V100-SXM3-32GB with 32 GB of RAM. Training on a single task took no more than one hour for any experiment.
@@ C.1 Standard Deviations Across Experimental Repetitions with Random Seeds
In the main paper, we observed methods in the continual setting performing well on the tasks they were trained on but completely failing on the unconfounded dataset in the disjoint case and also not reaching the joint accuracy in the strict case. To facilitate visual interpretation and focus on the main insights, we omitted reporting of standard deviations arising from experimental repetition with (five) random seeds in the main body. We now provide more details on why we consider these deviations to be negligible overall, but in particular with respect to the drawn conclusions.
@@
Other than these two methods, standard deviations are smaller than 1 percentage point of accuracy on all current tasks except for the naive training on the NN, where one outlier causes a higher deviation, the third task for DER with a still small standard deviation of 2.69, and for the cumulative setting, where it goes up to 3.65. For the latter, the higher deviation is probably due to the model being trained on more than the task 3 data at the same time, therefore not being able to prioritize it as much as other tasks, which almost reach perfect accuracy and thus have smaller deviations. We observe stronger deviations for the naive setup as well as EWC for old tasks. This fits with our observations of both approaches not successfully combating forgetting, resulting in partly random behavior on previous tasks. While there are some other slightly higher standard deviations of ER, DER, and BGS on old tasks as well, the largest deviation of these here is 5.79, which is still not relevant to any claims in our paper as the mean performances here are all below 60 and deviations as implied by the standard deviation still results in bad accuracies.
Most importantly, standard deviations on the unconfounded datasets are fairly small throughout. With the exception of the cumulative setting, we encounter the largest standard deviation for the NN with FROMP at around 2.77, with all others being smaller. This highlights the consistency of the failure to learn the ground truth rule. On the cumulative training setup, standard deviation on the third task moves up to 5.78 for NN (less than 2 for NeSy). Whereas this is, in fact, a large and relevant deviation from the mean, it is still a big difference from the more than 20 points difference that is observed by comparing the joint and the cumulative settings. This further emphasizes the NN’s preference for learning confounders over the ground truth in continual settings.
Table 4: ConCon Evaluation with Standard Deviations. Test accuracies on disjoint and strict for NN and NeSy, including standard deviations. Coloring is white until a standard deviation of 1 and then takes darker shades of purple until 10. With a few exceptions, most standard deviations are relatively small. Most importantly, performance on the unconfounded data is consistently bad.
@@
disjoint NN Naive Seq. 100.0$\pm$0.00 92.2$\pm$15.58 100.0$\pm$0.00 50.0$\pm$0.03 63.6$\pm$18.02 48.1$\pm$1.81
EWC [21] 100.0$\pm$0.00 99.9$\pm$0.07 100.0$\pm$0.00 50.1$\pm$0.21 60.6$\pm$13.97 47.7$\pm$2.18
FROMP [34] 100.0$\pm$0.00 99.8$\pm$0.14 99.4$\pm$1.00 99.2$\pm$1.53 97.9$\pm$2.41 50.2$\pm$0.17
ER [5] 100.0$\pm$0.00 99.6$\pm$0.36 100.0$\pm$0.03 100.0$\pm$0.00 99.6$\pm$0.22 50.1$\pm$0.10
DER [4] 100.0$\pm$0.00 99.9$\pm$0.07 100.0$\pm$0.03 99.2$\pm$1.31 98.1$\pm$1.47 50.2$\pm$0.08
BGS [25] 100.0$\pm$0.00 99.9$\pm$0.05 100.0$\pm$0.03 95.1$\pm$4.08 96.2$\pm$2.03 50.4$\pm$0.20
GDUMB [35] 64.1$\pm$18.38 62.0$\pm$9.70 69.5$\pm$9.51 63.6$\pm$10.71 61.2$\pm$8.93 48.8$\pm$0.76
@@
NeSy Naive Seq. 99.5$\pm$0.18 99.9$\pm$0.12 100.0$\pm$0.03 50.0$\pm$0.00 58.1$\pm$4.81 50.0$\pm$0.07
EWC [21] 99.6$\pm$0.28 100.0$\pm$0.03 99.9$\pm$0.05 50.2$\pm$0.35 67.2$\pm$9.12 50.2$\pm$0.18
FROMP [34] 99.6$\pm$0.31 98.9$\pm$0.29 98.4$\pm$0.69 97.9$\pm$1.45 97.3$\pm$0.72 50.5$\pm$0.23
ER [5] 99.5$\pm$0.17 99.8$\pm$0.05 99.9$\pm$0.04 99.2$\pm$0.19 99.3$\pm$0.23 50.7$\pm$0.18
DER [4] 99.6$\pm$0.22 99.7$\pm$0.13 99.9$\pm$0.08 98.6$\pm$0.64 96.9$\pm$1.12 50.9$\pm$0.05
BGS [25] 99.6$\pm$0.20 99.9$\pm$0.09 99.9$\pm$0.05 94.4$\pm$2.52 95.1$\pm$1.78 50.9$\pm$0.19
GDUMB [35] 99.4$\pm$0.12 97.8$\pm$0.89 98.5$\pm$0.72 98.0$\pm$0.86 98.2$\pm$0.90 50.3$\pm$0.15
@@
strict NN Naive Seq. 100.0$\pm$0.00 100.0$\pm$0.00 100.0$\pm$0.00 49.1$\pm$0.05 48.9$\pm$0.07 50.0$\pm$0.07
EWC [21] 100.0$\pm$0.00 99.9$\pm$0.10 100.0$\pm$0.05 49.1$\pm$0.05 48.9$\pm$0.08 50.0$\pm$0.07
FROMP [34] 100.0$\pm$0.00 59.9$\pm$8.49 69.2$\pm$4.92 67.8$\pm$10.71 59.1$\pm$6.35 51.1$\pm$2.77
ER [5] 100.0$\pm$0.00 97.7$\pm$0.50 96.4$\pm$0.59 54.6$\pm$1.76 59.6$\pm$2.71 51.0$\pm$1.24
DER [4] 100.0$\pm$0.00 98.7$\pm$0.92 96.5$\pm$2.69 52.8$\pm$4.13 55.5$\pm$5.79 49.1$\pm$0.26
BGS [25] 100.0$\pm$0.00 99.8$\pm$0.12 99.6$\pm$0.16 50.1$\pm$1.00 49.1$\pm$1.01 49.5$\pm$0.32
GDUMB [35] 92.5$\pm$2.77 62.7$\pm$8.34 53.6$\pm$2.32 57.5$\pm$5.20 54.2$\pm$2.89 50.4$\pm$0.50
@@
NeSy Naive Seq. 99.7$\pm$0.20 100.0$\pm$0.03 100.0$\pm$0.00 49.1$\pm$0.08 49.2$\pm$0.09 49.8$\pm$0.05
EWC [21] 99.9$\pm$0.08 99.9$\pm$0.07 100.0$\pm$0.03 49.1$\pm$0.08 49.4$\pm$0.49 50.1$\pm$0.19
FROMP [34] 99.8$\pm$0.12 83.4$\pm$9.01 74.7$\pm$7.18 85.1$\pm$7.65 76.3$\pm$10.56 56.2$\pm$1.17
ER [5] 99.8$\pm$0.17 99.8$\pm$0.15 98.8$\pm$0.11 68.0$\pm$1.42 83.6$\pm$0.71 61.2$\pm$0.82
DER [4] 99.7$\pm$0.23 99.4$\pm$0.15 99.5$\pm$0.24 58.8$\pm$1.30 73.2$\pm$3.91 56.5$\pm$1.05
BGS [25] 99.8$\pm$0.12 99.9$\pm$0.03 99.9$\pm$0.08 51.8$\pm$1.01 56.3$\pm$1.55 51.3$\pm$0.49
GDUMB [35] 99.4$\pm$0.24 85.6$\pm$1.39 82.7$\pm$1.64 80.1$\pm$2.42 84.2$\pm$2.33 56.7$\pm$1.09

### v3

@@ 5 Conclusion
We have introduced the first continually confounded dataset, ConCon, and have formally described its implications for systematic analysis. ConCon consists of two variants that implement different types of confounding. In our evaluations, we have empirically demonstrated that even perfect prevention of catastrophic forgetting is insufficient to avoid confounding in disjoint settings, where even joint training fails to generalize. Most importantly, we successfully confirmed the latter to not be true for our strict dataset, where joint training succeeds but continual cumulative training is veritably insufficient due to the manifestation of insidious continual confounding. This stands in stark contrast to the typical assumption that knowledge accumulation is the desirable solution in continual learning.
Our work thus provides the means to benchmark and further explore confounding in CL, for example, by investigating causal relationships (Mundt et al., 2023a) and related innovative new perspectives on, e.g., causal rehearsal (Churamani et al., 2023) or bias mitigation of spurious correlations (Lee et al., 2023). We envision that a deeper understanding of continual confounding can lead to overall better model performance in continual settings. Other modalities, such as text or time series, are not currently in the focus of ConCon and are left for future research.
Impact Statement
Our paper challenges the conventional view of cumulative learning as an upper bound when evaluating continual learning methods, thereby showing that such confounders have a detrimental effect on model performance even in what is conventionally considered a best-case scenario. The issue of confounders in machine learning represents a significant challenge. Making predictions based on such can not only lead to incorrect behavior but may also lead to unfair or discriminatory decisions (Schröder et al., 2023). In our specific case of ConCon, the dataset consists of synthetically generated geometric shapes and thus does not contain any personal or otherwise private information. At present, we do not envision any scenario in which it could be used with malicious intent. By means of ConCon, this paper aims to improve the evaluation of continual learning methods when confounding might be present, inspire new approaches to deal with these issues, and facilitate systematic investigation of such approaches in the future.
Acknowledgements
@@
    Training information for the NN (ResNet-18) and NeSy (Neuro-Symbolic concept learner) models.
@@
    Experimental results on the ConCon dataset using a generative model.
@@
    An experiment comparing joint and cumulative training on a dataset comprised of several ImageNet classes.
  • F:
    A statement on limitations of our paper and the introduced ConCon dataset.
  • G:
    A datasheet for datasets (Gebru et al., 2021) to document e.g., composition, collection process, and other aspects.
@@ Appendix A ConCon and the Code Framework
Along with our paper, we also publish the specific ConCon dataset [^5] https://zenodo.org/doi/10.5281/zenodo.10630481, as well as the data generation platform and the training code [^6] https://github.com/ml-research/concon to facilitate further research on continual confounding. We include a croissant (Akhtar et al., 2024) metadata record in our GitHub repository. [^7] https://github.com/ml-research/concon/blob/main/croissant_concon.json ConCon uses the BSD 2-Clause ”Simplified” License, and our code for generating the data and running the experiments is released under the MIT license.
@@ A.1 Dataset Details
Recall that ConCon is built on top of CLEVR (Johnson et al., 2017) and thus consists of images representing geometric objects with different properties. Each image is an RGB image of size 224$\times$224 depicting four objects. In our dataset, an object can be described by four properties, namely shape (cube, sphere, cylinder), material (metal or rubber), size (small or large), and color (gray, red, blue, green, brown, purple, cyan, yellow). Data is generated with respect to the ground truth rule and confounders as described in the main paper. Relationships between objects are not used in ConCon. We show several samples in Figure 4.
As we describe in detail in our main paper, the positive instances in the dataset fulfill the ground truth rule ($g$) and the negative instances do not ($\neg g$). In addition, it is possible that a confounding feature must apply ($c_{t}$) or must not apply ($\neg c_{t}$), depending on the task (denoted by subscript $t$) and the dataset variant (disjoint or strict). The ground truth $g$ is defined as sphere and small cube. The confounders, in the order of tasks, are $c_{1}$: blue, $c_{2}$: metal, and $c_{3}$: large. We summarize the rules that should hold for respective tasks in Tables 3 (disjoint) and 4 (strict).
@@
Table 4: Rules strict: a confounding feature must appear in the positive images but may not appear in the negative images of its respective task. The confounding features may, but are not required to, appear in images of other tasks.
@@ A.2 Code Framework
It is possible to generate further instances or make changes to the data generation process by using our code repository. The logic in the form of ground truth and confounders, which determines the data generation, can be set in separate files. Note that the intended behavior for our data generation consists of using simple confounders that are only required to be satisfied on a single object. Here, our code could easily be adapted, for example, to support the generation of images based on multi-object confounders by changing the selection of object properties in the respective function. In a similar spirit, both CLEVR and ConCon are generated with the Blender graphics simulation software, allowing for the inclusion of entirely new attributes and confounding features themselves. As such, extensions to, e.g., novel color choices are realizable with either trivial or modest amounts of specification/change in code, or other already available properties (such as relationships) could easily be leveraged in future dataset creation and empirical investigation. Whereas these even more challenging setups may certainly be of interest in the long term, we note that our presently chosen configuration already highlights the catastrophic behavior of investigated methods and exposes insidious continual confounding, making solving our generated ConCon variant first desirable.
@@ Appendix B Experimental Details
Methods. Continual learning methods aim to alleviate forgetting when training and updating models continually. We here give additional information on the CL methods that we consider in our evaluation in the main paper: i) EWC (Kirkpatrick et al., 2017) is a method that regularizes weights relevant to the previous task, ii) FROMP (Pan et al., 2020) stores examples relevant to the previous task and performs functional regularization while training on the current tasks, iii) ER (Chaudhry et al., 2019) maintains a fixed-size memory buffer of random samples from prior tasks and interleaves them while training on the current task, iv) DER (Buzzega et al., 2020) also maintains a memory buffer that stores past samples and predictions, and in addition, the loss function includes a cross-entropy loss for the current task and a regularization term using the stored logits for past tasks, v) BGS (Lee et al., 2024) tackles the problem of spurious correlations in continual learning by storing samples that are balanced across the class labels and spurious features, vi) GDUMB (Prabhu et al., 2020) stores uniformly balanced past samples but is not a conventional continual learner as the model is trained from scratch at every new task, and vii) PNNs (Rusu et al., 2022) progressively grow the network by adding new subnetworks for each task, expanding the structure and parameters of previously learned networks.
@@
NN Training. We use a ResNet-18 (He et al., 2016) model, a convolutional neural network-based architecture with skip connections. Following standard deep learning practice, the model is trained with an Adam optimizer (Kingma & Ba, 2014) with a mini-batch size of 64 and a learning rate of 0.001. The model was trained for 50 epochs for all methods but FROMP, where we trained the NN model for 30 epochs, as the performance became worse when training for 50.
@@
NeSy Training. We use a neuro-symbolic concept learner (NeSy) (Stammer et al., 2021) that is equipped with a Slot attention model and a Set Transformer. The Slot-Attention module of NeSy is pre-trained on the original CLEVR dataset (Locatello et al., 2020) for the purpose of extracting accurate neuro-symbolic intermediate representations (e.g., disentangled slots for color and shape). We use the (perfectly accurate as a consequence of simulation) structured information from the embeddings of the attention module to then learn only the Set Transformer for the ConCon dataset. The latter is trained using an Adam optimizer (Kingma & Ba, 2014) with a mini-batch size of 64 and a learning rate of 0.001. The model was trained for 150 epochs. We apply transformations to resize the images to 128$\times$128 to fit the constraints set by prior works (Stammer et al., 2021).
@@
Hyperparameter Tuning. We did not excessively tune hyperparameters as all models achieved good solutions when expected in their baseline scenario. As explained in the main body and discussed in much more detail in the next section, observed failure cases are systematic and structural in nature - a consequence of the continual learning set-up rather than hyperparameter values.
@@
Technical Details. We created the dataset using an AMD EPYC 7742 64-Core Processor CPU. All models were built using PyTorch and trained and evaluated on an NVIDIA Tesla V100-SXM3-32GB with 32 GB of RAM. Training on a single task took no more than one hour for any experiment, with the exception of PNNs, which sometimes exceeded one hour.
@@ C.1 Standard Deviations Across Experimental Repetitions with Random Seeds
In the main paper, we observed methods in the continual setting performing well on the tasks they were trained on but completely failing on the unconfounded dataset in the disjoint case and also not reaching the joint accuracy in the strict case. To facilitate visual interpretation and focus on the main insights, we omitted reporting of standard deviations arising from experimental repetition with five random seeds in the main body. We now provide more details on why we consider these deviations to be negligible overall, but in particular with respect to the drawn conclusions.
@@
Other than these two methods, standard deviations are smaller than 1 percentage point of accuracy on all current tasks except for the naive training on the NN, where one outlier causes a higher deviation, the third task for DER with a still small standard deviation of 2.69, and for the shuffled and cumulative settings, where it goes up to 3.85. For the latter two, the higher deviation is probably due to the model being trained on equal amounts of data from different tasks at the same time, therefore not being able to prioritize single tasks as much as other tasks, which almost reach perfect accuracy and thus have smaller deviations. We observe stronger deviations for the naive setup as well as EWC for old tasks. This fits with our observations of both approaches not successfully combating forgetting, resulting in partly random behavior on previous tasks. While there are some other slightly higher standard deviations of ER, DER, and BGS on old tasks as well, the largest deviation of these here is 5.79, which is still not relevant to any claims in our paper as the mean performances here are all below 60 and deviations as implied by the standard deviation still results in bad accuracies.
Most importantly, standard deviations on the unconfounded datasets are fairly small throughout. With the exception of the shuffled and cumulative settings, we encounter the largest standard deviation for the NN with FROMP at around 2.77, with all others being smaller. This highlights the consistency of the failure to learn the ground truth rule. On the shuffled training setup, standard deviation reaches 4.63 (less than 1 for NeSy) and on the cumulative training setup, standard deviation on the third task moves up to 5.78 for NN (less than 1 for NeSy). Whereas this is, in fact, a large and relevant deviation from the mean, it is still a big difference from the more than 20 points difference that is observed by comparing the joint and the cumulative settings or the more than 15 points difference between shuffled and cumulative. This further emphasizes the NN’s preference for learning confounders over the ground truth in continual settings.
Table 5: ConCon Evaluation with Standard Deviations. Test accuracies on disjoint and strict for NN and NeSy, including standard deviations. Coloring is white until a standard deviation of 1, and then takes darker shades of purple until 10. With a few exceptions, most standard deviations are relatively small. Most importantly, performance on the unconfounded data is consistently bad.
@@
disjoint NN Naive Sequential 100.0$\pm$0.00 92.2$\pm$15.58 100.0$\pm$0.00 50.0$\pm$0.03 63.6$\pm$18.02 48.1$\pm$1.81
EWC (Kirkpatrick et al., 2017) 100.0$\pm$0.00 99.9$\pm$0.07 100.0$\pm$0.00 50.1$\pm$0.21 60.6$\pm$13.97 47.7$\pm$2.18
FROMP (Pan et al., 2020) 100.0$\pm$0.00 99.8$\pm$0.14 99.4$\pm$1.00 99.2$\pm$1.53 97.9$\pm$2.41 50.2$\pm$0.17
ER (Chaudhry et al., 2019) 100.0$\pm$0.00 99.6$\pm$0.36 100.0$\pm$0.03 100.0$\pm$0.00 99.6$\pm$0.22 50.1$\pm$0.10
DER (Buzzega et al., 2020) 100.0$\pm$0.00 99.9$\pm$0.07 100.0$\pm$0.03 99.2$\pm$1.31 98.1$\pm$1.47 50.2$\pm$0.08
BGS (Lee et al., 2024) 100.0$\pm$0.00 99.9$\pm$0.05 100.0$\pm$0.03 95.1$\pm$4.08 96.2$\pm$2.03 50.4$\pm$0.20
GDUMB (Prabhu et al., 2020) 64.1$\pm$18.38 62.0$\pm$9.70 69.5$\pm$9.51 63.6$\pm$10.71 61.2$\pm$8.93 48.8$\pm$0.76
PNNs (Rusu et al., 2022) 100.0$\pm$0.00 99.8$\pm$0.05 100.0$\pm$0.00 100.0$\pm$0.00 99.8$\pm$0.05 47.9$\pm$2.36
Shuffled 100.0$\pm$0.03 99.9$\pm$0.03 100.0$\pm$0.03 100.0$\pm$0.03 99.8$\pm$0.05 50.0$\pm$0.03
@@
NeSy Naive Sequential 99.5$\pm$0.18 99.9$\pm$0.12 100.0$\pm$0.03 50.0$\pm$0.00 58.1$\pm$4.81 50.0$\pm$0.07
EWC (Kirkpatrick et al., 2017) 99.6$\pm$0.28 100.0$\pm$0.03 99.9$\pm$0.05 50.2$\pm$0.35 67.2$\pm$9.12 50.2$\pm$0.18
FROMP (Pan et al., 2020) 99.6$\pm$0.31 98.9$\pm$0.29 98.4$\pm$0.69 97.9$\pm$1.45 97.3$\pm$0.72 50.5$\pm$0.23
ER (Chaudhry et al., 2019) 99.5$\pm$0.17 99.8$\pm$0.05 99.9$\pm$0.04 99.2$\pm$0.19 99.3$\pm$0.23 50.7$\pm$0.18
DER (Buzzega et al., 2020) 99.6$\pm$0.22 99.7$\pm$0.13 99.9$\pm$0.08 98.6$\pm$0.64 96.9$\pm$1.12 50.9$\pm$0.05
BGS (Lee et al., 2024) 99.6$\pm$0.20 99.9$\pm$0.09 99.9$\pm$0.05 94.4$\pm$2.52 95.1$\pm$1.78 50.9$\pm$0.19
GDUMB (Prabhu et al., 2020) 99.4$\pm$0.12 97.8$\pm$0.89 98.5$\pm$0.72 98.0$\pm$0.86 98.2$\pm$0.90 50.3$\pm$0.15
PNNs (Rusu et al., 2022) 99.5$\pm$0.25 100.0$\pm$0.00 100.0$\pm$0.00 99.6$\pm$0.27 100.0$\pm$0.00 49.6$\pm$0.23
Shuffled 99.8$\pm$0.09 99.7$\pm$0.09 99.9$\pm$0.05 99.7$\pm$0.13 99.9$\pm$0.03 50.8$\pm$0.07
@@
strict NN Naive Sequential 100.0$\pm$0.00 100.0$\pm$0.00 100.0$\pm$0.00 49.1$\pm$0.05 48.9$\pm$0.07 50.0$\pm$0.07
EWC (Kirkpatrick et al., 2017) 100.0$\pm$0.00 99.9$\pm$0.10 100.0$\pm$0.05 49.1$\pm$0.05 48.9$\pm$0.08 50.0$\pm$0.07
FROMP (Pan et al., 2020) 100.0$\pm$0.00 59.9$\pm$8.49 69.2$\pm$4.92 67.8$\pm$10.71 59.1$\pm$6.35 51.1$\pm$2.77
ER (Chaudhry et al., 2019) 100.0$\pm$0.00 97.7$\pm$0.50 96.4$\pm$0.59 54.6$\pm$1.76 59.6$\pm$2.71 51.0$\pm$1.24
DER (Buzzega et al., 2020) 100.0$\pm$0.00 98.7$\pm$0.92 96.5$\pm$2.69 52.8$\pm$4.13 55.5$\pm$5.79 49.1$\pm$0.26
BGS (Lee et al., 2024) 100.0$\pm$0.00 99.8$\pm$0.12 99.6$\pm$0.16 50.1$\pm$1.00 49.1$\pm$1.01 49.5$\pm$0.32
GDUMB (Prabhu et al., 2020) 92.5$\pm$2.77 62.7$\pm$8.34 53.6$\pm$2.32 57.5$\pm$5.20 54.2$\pm$2.89 50.4$\pm$0.50
PNNs (Rusu et al., 2022) 100.0$\pm$0.00 99.8$\pm$0.09 99.8$\pm$0.42 100.0$\pm$0.00 99.8$\pm$0.09 49.2$\pm$1.23
Shuffled 79.9$\pm$3.11 93.4$\pm$3.85 95.3$\pm$2.14 95.2$\pm$3.21 96.7$\pm$1.76 89.5$\pm$4.63
@@
NeSy Naive Sequential 99.7$\pm$0.20 100.0$\pm$0.03 100.0$\pm$0.00 49.1$\pm$0.08 49.2$\pm$0.09 49.8$\pm$0.05
EWC (Kirkpatrick et al., 2017) 99.9$\pm$0.08 99.9$\pm$0.07 100.0$\pm$0.03 49.1$\pm$0.08 49.4$\pm$0.49 50.1$\pm$0.19
FROMP (Pan et al., 2020) 99.8$\pm$0.12 83.4$\pm$9.01 74.7$\pm$7.18 85.1$\pm$7.65 76.3$\pm$10.56 56.2$\pm$1.17
ER (Chaudhry et al., 2019) 99.8$\pm$0.17 99.8$\pm$0.15 98.8$\pm$0.11 68.0$\pm$1.42 83.6$\pm$0.71 61.2$\pm$0.82
DER (Buzzega et al., 2020) 99.7$\pm$0.23 99.4$\pm$0.15 99.5$\pm$0.24 58.8$\pm$1.30 73.2$\pm$3.91 56.5$\pm$1.05
BGS (Lee et al., 2024) 99.8$\pm$0.12 99.9$\pm$0.03 99.9$\pm$0.08 51.8$\pm$1.01 56.3$\pm$1.55 51.3$\pm$0.49
GDUMB (Prabhu et al., 2020) 99.4$\pm$0.24 85.6$\pm$1.39 82.7$\pm$1.64 80.1$\pm$2.42 84.2$\pm$2.33 56.7$\pm$1.09
PNNs (Rusu et al., 2022) 99.8$\pm$0.07 100.0$\pm$0.03 100.0$\pm$0.00 99.8$\pm$0.09 100.0$\pm$0.03 49.9$\pm$0.05
Shuffled 87.8$\pm$0.66 92.5$\pm$0.59 89.2$\pm$0.60 88.2$\pm$0.35 92.6$\pm$0.79 70.1$\pm$0.51